### v2

# Improving Low-Resource Sequence Labeling with Knowledge Fusion and Contextual Label Explanations

Abstract
Sequence labeling remains a significant challenge in low-resource, domain-specific scenarios, particularly for character-dense languages. Existing methods primarily focus on enhancing model comprehension and improving data diversity to boost performance. However, these approaches still struggle with inadequate model applicability and semantic distribution biases in domain-specific contexts. To overcome these limitations, we propose a novel framework that combines an LLM-based knowledge enhancement workflow with a span-based Knowledge Fusion for Rich and Efficient Extraction (KnowFREE) model. Our workflow employs explanation prompts to generate precise contextual interpretations of target entities, effectively mitigating semantic biases and enriching the model’s contextual understanding. The KnowFREE model further integrates extension label features, enabling efficient nested entity extraction without relying on external knowledge during inference. Experiments on multiple Chinese domain-specific sequence labeling datasets demonstrate that our approach achieves state-of-the-art performance, effectively addressing the challenges posed by low-resource settings.
[…]
Improving Low-Resource Sequence Labeling with Knowledge Fusion and Contextual Label Explanations
Peichao Lai1, Jiaxin Gan2, Feiyang Ye2, Wentao Zhang1, Fangcheng Fu1, Yilei Wang2, Bin Cui1,* 1School of Computer Science, Peking University, 2College of Computer and Data Science, Fuzhou University *bin.cui@pku.edu.cn
1 Introduction
Sequence labeling is a fine-grained information extraction (IE) task that includes sub-tasks such as named entity recognition (NER) and part-of-speech (POS) tagging, playing a critical role in various downstream natural language processing (NLP) applications. However, in low-resource scenarios, sequence labeling for character-dense languages like Chinese presents unique challenges. The absence of explicit word boundaries makes accurate label inference more complex, while the scarcity of domain-specific data further limits the model’s ability to effectively capture label distributions.
Previous studies on enhancing sequence labeling in low-resource with character-dense languages can be broadly classified into two streams: (1) Model-Centric Optimization. These methods focus on enhancing model’s comprehension to detect implicit word boundaries and contextual signals through feature engineering. For instance, lexical features are injected via lexicon matching networks (Zhang and Yang, 2018a; Li et al., 2020; Liu et al., 2021; Wu et al., 2021) or prompt templates (Ma et al., 2022b; Shen et al., 2023; Chen et al., 2021b) to strengthen entity boundary detection, while others employ knowledge transfer mechanisms like Gaussian embeddings (Si et al., 2024; Das et al., 2022), prompt-based metrics (Chen et al., 2023; Lai et al., 2022) and contrastive learning (Huang et al., 2022; Zhang et al., 2024) to distill knowledge into target domains. (2) Data-Centric Augmentation. Meanwhile, other studies concentrate on using data augmentation through altering entity label information (Hu et al., 2023; Yang et al., 2018) and extracting knowledge from the external environment (Cai et al., 2023; Chen et al., 2021a; Yaseen and Langer, 2021) to enrich the dataset. With the advent of large language models (LLMs), recent findings leverage their generative capabilities to enhance the diversity of entity and sentence synthesis (Kang et al., 2024; Ye et al., 2024).
However, critical limitations persist when applying these solutions to specialized domains: (1) Limited Model Applicability. Existing model-centric approaches for character-dense languages often struggle to integrate diverse feature types and labels effectively, restricting the flexibility and richness of feature injection. Moreover, they face challenges in handling nested entities, further constraining their adaptability. Additionally, most approaches depend on rigid feature injection frameworks and complex input configurations to enhance word boundary detection, leading to increased reliance on supplementary structures during inference and higher usage costs. (2) Variability in Label Distribution. Existing data-centric augmentation techniques suffer from domain distribution bias. Differences in entity type definitions and sample semantics across domains lead to misaligned label priors and entity contexts, significantly compromising both the quality of augmented data and zero-shot generalization. As illustrated in Figure 1, the aforementioned challenges of structural inflexibility and distribution discrepancy collectively constrain current methods’ practical efficacy, motivating our investigation into a unified framework that simultaneously resolves model architecture limitations and domain adaptation barriers.
Figure 1: Distinctions between our method and existing methods in terms of model and data.
In this task, we adopt two key strategies for improving low-resource sequence labeling in character-dense languages: (i) enhancing the utilization of non-entity features and (ii) improving the model’s contextual understanding of target entities. To achieve these objectives, we propose a novel LLM-based data augmentation framework. Our approach begins by designing extraction prompts to identify and extract potential non-target entity features from the input text, thereby maximizing the utilization of non-entity features. To address the issue of limited model applicability, we introduce a span-based model, Knowledge Fusion for Rich and Efficient Extraction (KnowFREE) model, which supports nested label annotations and integrates extension label features by a local multi-head attention module. Unlike prior methods, KnowFREE learns rich contextual information during training without relying on external knowledge at inference. To tackle the issue of variability in label distribution, we incorporate explanation prompts inspired by label explanation techniques (Golde et al., 2024; Yang and Katiyar, 2020; Ma et al., 2022a), generating precise explanations for target entities in their contexts. This step further strengthens the model’s contextual understanding of target entities. This novel use of LLMs for label interpretation synthesis effectively mitigates semantic distribution bias, outperforming traditional data augmentation techniques in low-resource scenarios. We evaluate our framework on multiple Chinese domain-specific sequence labeling datasets, and experimental results demonstrate its effectiveness in addressing the challenges of low-resource settings.
The contributions of our work can be summarized as follows:
(1) New method. We propose a span-based KnowFREE model that supports nested label annotations and integrates multi-label features by a local multi-head attention module, which can be used without relying on external knowledge at inference.
(2) New perspective. To the best of our knowledge, we are the first to propose an approach that supports the seamless integration of extension label features within the model while eliminating the need for external features during inference.
(3) State-of-the-art performance. Experimental results demonstrate that our approach achieves outstanding performance on low-resource sequence labeling tasks.
2 Related Work
Span-based Sequence Labeling Methods Span-based sequence labeling methods have gained prominence for their ability to address overlapping and nested entities effectively. Early works, such as Dozat and Manning (2017); Yu et al. (2020) introduced biaffine models to capture sentence-wide structures and score span boundaries for accurate information extraction. Based on this, Su et al. (2022) proposed the Global Pointer model, optimizing the biaffine transformation’s weight matrix and bias terms to boost efficiency and precision in span-based NER. In parallel, Shen et al. (2022) introduced a parallel instance query network for simultaneous entity extraction. Then, (Yan et al., 2023) proposed a multi-head biaffine mechanism combined with CNNs to capture local span features, achieving improved performance, while (Li et al., 2022) combined bilinear classifiers with dilated convolutions post-CLN to refine span-level relation classification. For few-shot NER, (Wang et al., 2022) introduced SpanProto, which integrates prototype-based classification with a contrastive loss to effectively separate non-target spans from prototype clusters, demonstrating strong performance in low-resource scenarios.
Sequence Labeling via LLMs The advent of LLMs (OpenAI, 2023; Bai et al., 2023; Touvron et al., 2023) has introduced innovative strategies for sequence labeling. Generative approaches utilizing in-context learning (ICL) allow LLMs to perform sequence labeling tasks directly (Jiang et al., 2024). Leveraging their extensive knowledge and flexible template structures, LLMs have also shown promise in supporting cross-domain entity recognition (Nandi and Agrawal, 2024) and tackling nested entity challenges (Kim et al., 2024). In few-shot scenarios, methods like ProgGen (Heng et al., 2024) exploit LLMs to iteratively synthesize attribute-rich training data by incorporating domain-specific features. Chen et al. (2024) integrates a small span detection model with LLM validation to enhance span selection and address domain-specific challenges.
3 Method
In this section, we will introduce the knowledge enhancement workflow and the specific structure of our KnowFree model for information extraction. The overall framework is shown in Figure 2.
Figure 2: The workflow of knowledge enhancement.
3.1 Workflow of Knowledge Enhancement
In the knowledge enhancement workflow, we leverage LLMs to annotate potential entity information in the source sample and provide additional descriptions of entities. This enhances the utilization of non-entity features and improves the model’s comprehension of the context in which the target entity appears. Our workflow consists of two main pipelines, which we describe in detail below.
Label Extension Annotation. In low-resource scenarios, non-entity segments in the sentence may contain additional non-target entity features, while data samples with flat-only entities may potentially contain nested entity information around the target entities. Leveraging these potential feature information can enhance the model’s capacity to comprehend the fine-grained semantics and the ability to distinguish entity boundaries in character-dense languages. To achieve this, we utilize the LLM’s general knowledge to generate extension entity tags, word segmentation tags, and part-of-speech (POS) tags for the source samples.
Formally, we can denote the source samples as $S=\{s_{1},s_{2},\dots,s_{n}\}$, and the LLM as $L$, where $n$ is the number of samples. We then constructed a prompt for each type of tag extraction, denoted as $\mathcal{P}_{ent}$, $\mathcal{P}_{seg}$, and $\mathcal{P}_{pos}$, correspondingly. The extension tags set can be computed as:
$\hat{E}^{\left(k\right)}=\bigcup_{i=1}^{n}\hat{E}^{\left(k\right)}_{i},\quad\hat{E}^{\left(k\right)}_{i}=L(s_{i},\mathcal{P}_{k}),$ (1)
where $\hat{E}^{\left(k\right)}_{i}$ represents the extension set using prompt $\mathcal{P}_{k}$ extracted from sentence $s_{i}$.
We assume that accurate and diverse extension tags can significantly improve the model’s comprehension of context and its capacity for entity detection. Then, there are several issues that need to be solved in the results of extraction. (1) The uncertainty generation by LLMs leads to the presence of tags in the extension entity set in various textual formats yet denotes the same label category. Directly introducing these labels will significantly disrupt the model’s assessment of the entity type. (2) Word segmentation usually produces boundary labels with low information entropy. However, POS tagging based on LLMs sometimes has missing annotations. Notably, POS tagging labels inherently include implicit word boundary information. Therefore, combining the outputs of these two processes is expected to yield better results.
To address the first issue, we use LLM to generate synonymous label mapping, and combine entity clustering algorithm to achieve synonymous label merging. Let denote the extension entity set as $\hat{E}^{\mathrm{ent}}=\{(e_{i},t_{i})\mid e_{i}\in\mathcal{E},t_{i}\in\mathcal{T}\}$, where $\mathcal{E}$ is the set of entities and $\mathcal{T}$ is the set of entity types. We compute the synonymous tag set by using the synonymous tag merge prompt $\mathcal{P}_{merge}$:
$\mathcal{T}_{s}=L(\mathcal{T},\mathcal{P}_{\mathrm{merge}}),$ (2)
$\mathcal{T}_{s}=\{\tilde{T}_{i}:\{T_{i1},T_{i2},\ldots,T_{ir}\}\mid i\in[1,m]\},$ (3)
where $\tilde{T}_{i}$ is the standard label, $T_{ij}$ is the synonymous label, and $m$, $r$ is the number of standard label and its corresponding synonymous labels, respectively. This method is determined by the LLM’s literal interpretation of the label, which may not accurately align the semantic spatial distribution of entities in the target domain, and therefore is not completely reliable. We further compute the vector representation of each entity-label pair by using a sentence embedding model $\mathcal{M}$, and vector set of $T_{k}$ can be represented as:
$\mathbf{V}_{k}=\{\mathbf{v}_{i}\mid(e_{i},t_{i})\in\hat{E}^{\text{ent}},\mathbf{t}_{i}\in T_{k}\},$ (4)
$\mathbf{v}_{i}=\mathcal{M}(x_{i}),$ (5)
where the input $x_{i}$ is the concatenation of $e_{i}$ and $t_{i}$ with the template of “[$e_{i}$] is [$t_{i}$]”. The center point and mean radius of the vector set for $T_{k}$ can be calculated as:
$\mathbf{c}_{k}=\frac{1}{|\mathbf{V}_{k}|}\sum_{\mathbf{v}_{i}\in\mathbf{V}_{k}}\mathbf{v}_{i},$ (6)
$r_{k}=\frac{1}{p}\sum_{j=1}^{p}\|\mathbf{v}_{j}-\mathbf{c}_{k}\|,$ (7)
where $p$ represents Top-$p$ samples with the largest distance from $\mathbf{c}_{k}$. Then, for each standard label $\tilde{T}_{i}$, we identify the synonymous label vector set $\mathbf{V}_{i,max}$ that contains the largest number of samples and has the maximum radius, designating it as the reference vector set. Subsequently, we evaluate whether other synonymous label vector sets $\mathbf{V}_{i,j}$ satisfy the following condition:
$\|\mathbf{c}_{j}-\mathbf{c}_{max}\|\leq\epsilon\cdot r_{max}.$ (8)
If the condition holds, $T_{j}$ is merged into $\tilde{T}_{i}$; Otherwise, $T_{j}$ is treated as an independent standard label. To address the second issue, we initially compute the word segmentation set $\hat{E}^{\mathrm{seg}}$ utilizing $\mathcal{P}_{seg}$, then we employ $\hat{E}^{\mathrm{seg}}$, $\mathcal{P}_{pos}$ and the source sample as input for the LLM to generate the part of speech without omission. This effectively ensures that all words in the sample are tagged with part of speech, and also improves the diversity of word segmentation features.
Finally, we merge all the extension entities with standardized labels into the original data to obtain the fusion samples. We use our KnowFree model to first train an annotation model on the fusion samples, which can be used for entity annotation in the next pipeline.
Enriched Explanation Synthesis. Injecting extension entity features has been proven to enhance model performance in many cases, its impact remains constrained in the following situations: (1) The datasets with short sentence contains a high proportion of target entities, leading to a relatively low amount of non-entity information in the sample. This results in a diminished validity of the external features introduced. (2) As the quantity of samples diminishes, their scarcity will emerge as the principal factor limiting performance enhancement. Enhancing the exploitation of non-entity features in the sample yields marginal performance enhancement. Both instances highlight the necessity of expanding the sample size. Nonetheless, the divergence in semantic distribution indicates that incorporating samples from outside the source domain, along with directly employing LLM to generate new sentences from existing entities, may introduce potential noise that could significantly affect model performance.
To tackle the aforementioned challenges, we employ LLM to generate entity explanations within the current samples, which aims to leverage the knowledge embedded within the LLM and the connections between samples and entities to produce the precise meaning of target entities within their context. This approach can mitigate noise caused by semantic distribution shifts in synthetic samples. Specifically, we define two types of explanation prompts: the Entity Explanation Prompt $\mathcal{P}_{exp}$ and the Extension Description Prompt $\mathcal{P}_{ext}$ for samples with and without target entities, respectively. The function of $\mathcal{P}_{exp}$ takes the source sample and its corresponding target entity as input, aiming to generate a detailed explanation of the entity’s contextual meaning. Meanwhile, $\mathcal{P}_{ext}$ focuses on extracting and explaining key phrases from the text, with only the source sample as input. Additionally, to enrich the semantic representation of the source data, both prompts instruct the LLM to act the role of a domain expert, providing accessible and detailed explanations of the target entities to a hypothetical audience. This strategy encourages the LLM to generate comprehensive, contextually relevant, and easy-to-understand explanations, enhancing the overall semantic clarity of the dataset.
Next, we generate enriched explanations using the explanation prompts and proceed with annotating the entities in these synthetic texts through a two-branch pipeline: one for target entity annotation and the other for extension entity extraction. The frozen KnowFree model, trained in the previous pipeline, is used to annotate the target entities. For extracting extension entities, we apply the same method as in the previous pipeline. Finally, all entities are integrated into the enriched explanations to produce synthetic samples. As shown in Figure 2, we can combine fusion samples and synthetic samples and retrain on the KnowFree model to enhance its performance in the few-shot scenario.
3.2 Structure of KnowFree Model
Figure 3: The structure of KnowFree model.
To support the fusion of multi-label knowledge, our proposed model is implemented based on the biaffine structure. The structure of our KnowFree model is illustrated in Figure 3. Compared to previous approaches that involved external feature injection, our approach avoids the creation of additional feature injection modules at the input stage. Both extension entity labels and target entity labels are input together into the model for training, with their weights adjusted through hyperparameters during the loss function computation. During inference, we mask the weights of the extension entity labels, ensuring that the model only predicts target entities. This approach significantly simplifies the model while improving its overall efficiency. We additionally incorporate a local multi-head attention layer into the model to improve the interaction of multi-label features among neighborhoods during the training process.
Formally, let denote $X,\hat{X}$ as inputs of the fusion sample and the synthetic sample, respectively. The target entity spans of each sample is represented as $[s_{i},e_{i},l_{i}]$, where $s_{i},e_{i}$ are the start and end indices of the entity span, and $l_{i}$ is the label type. The extension entity span is represented as $[s_{j},e_{j},\tilde{l}_{j}]$, where $\tilde{l}_{j}$ is the extension label type. We adopt a pre-trained encoder to compute the hidden states $H\in\mathbb{R}^{L\times D}$ for each input, where $L$ is the length of the input sequence, $D$ is the dimension of hidden states. We then compute the encoding of the entity’s start and end positions:
$H_{s}=\sigma(H\mathbf{W}_{s}),\quad H_{e}=\sigma(H\mathbf{W}_{e}),$ (9)
where $\mathbf{W}_{s},\mathbf{W}_{e}\in\mathbb{R}^{D\times D^{\prime}}$ are learnable weight matrics, $D^{\prime}$ is the feature hidden size, and $\sigma$ is the activation function. The biaffine matrix of spans is computed by a multi-head Biaffine decoder:
$H_{\text{B}}=\mathcal{F}_{\text{MHB}}(H_{s},H_{e}),$ (10)
where $\mathcal{F}_{\text{MHB}}$ represents the multi-head Biaffine decoder, and $H_{\text{B}}\in\mathbb{R}^{L\times L\times\tilde{D}}$, $\tilde{D}$ is the hidden size of the biaffine matrix.
To improve the interactivity between multi-label features and span neighborhoods, we introduce a local multi-head attention layer to generate a mask for local multi-head attention, ensuring that each token can only interact with its neighboring tokens:
$M[i,j]=\begin{cases}0&\text{if }|i-j|\leq w,\\ -\infty&\text{otherwise}.\end{cases}$ (11)
where $M\in\mathbb{R}^{L\times L}$, $\omega$ is the window size of local multi-head attention. The input feature $H_{\text{B}}$ is divided into $K$ attention heads, each with a reduced feature dimension of $\tilde{D}/K$. For each attention head, we compute the query $Q$, key $K$, and value $V$ matrices via linear projections of the input:
$Q^{(k)}=H_{\text{B}}\mathbf{W}_{Q}^{(k)},K^{(k)}=H_{\text{B}}\mathbf{W}_{K}^{(k)},V^{(k)}=H_{\text{B}}\mathbf{W}_{V}^{(k)},$
where $\mathbf{W}_{Q}^{(k)},\mathbf{W}_{K}^{(k)},\mathbf{W}_{V}^{(k)}\in\mathbb{R}^{\tilde{D}\times(\tilde{D}/{K})}$ are learnable projection matrices for the $k$-th head. We then compute the scaled dot-product attention for each attention head:
$\mathcal{A}(Q^{(k)},K^{(k)},V^{(k)},M)=\text{Softmax}\left(\frac{Q^{(k)}(K^{(k)})^{T}+M}{\sqrt{d_{k}}}\right)V^{(k)},$
where $d_{k}=\tilde{D}/{K}$ is the dimension size of the $k$-th head. The outputs from all attention heads are concatenated as:
$\mathcal{\tilde{A}}=[\mathcal{A}_{1};\mathcal{A}_{2};\ldots;\mathcal{A}_{K}],$ (14)
$H_{\text{attn}}=\text{LayerNorm}(\mathcal{\tilde{A}}),$ (15)
where $\mathcal{\tilde{A}}\in\mathbb{R}^{L\times L\times\tilde{D}}$, and the output feature is finally computed with a layer normalization. To enhance training stability and retain the initial biaffine features, we apply a residual connection followed by layer normalization:
$H_{\text{G}}=\text{LayerNorm}(H_{\text{B}}+H_{\text{attn}}).$ (16)
We then use the fully connection layer to map the sum of $H_{\text{G}}$ and $H_{\text{B}}$ into the number of entity tags:
$P=\sigma^{\ast}(\mathbf{W}_{O}(H_{\text{B}}+H_{\text{G}})+\mathbf{b}),$ (17)
where $\mathbf{W}_{O}\in\mathbb{R}^{\tilde{D}\times(N_{\text{tgt}}+N_{\text{ext}})}$ is the learnable weight matrix, $\sigma^{\ast}$ is the activation function, and $\mathbf{b}\in\mathbb{R}^{(N_{\text{tgt}}+N_{\text{ext}})}$ is the bias. $N_{\text{tgt}},N_{\text{ext}}$ are the number of target entity label and extension entity label, respectively. The binary cross entropy is employed to compute the loss. To prevent the model from overly concentrating on features of extension entities, the loss function is defined as:
$\mathcal{L}=-(\sum_{\begin{subarray}{c}0\leq i,\\ j<N_{\text{tgt}}\end{subarray}}y_{i,j}\log P_{i,j}+\alpha\sum_{\begin{subarray}{c}N_{\text{tgt}}\leq i,\\ j<N_{\text{ext}}\end{subarray}}y_{i,j}\log P_{i,j}),$ (18)
where $y$ refers to the ground truth labels, $\alpha$ is the weight parameter. Similarly, we control the influence of the quantity and noise in synthetic samples with another weight parameter. The final loss is a weighted sum of the original and synthetic losses:
$\mathcal{L}_{\text{final}}=\mathcal{L}+\beta\mathcal{L}_{syn}.$ (19)
4 Experiment
Dataset / Model Weibo Youku Taobao Resume
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003) 56.57 60.91 66.52 68.02 70.57 74.92 68.78 71.88 74.74 90.19 92.35 93.43
FLAT (Li et al., 2020) 57.75 59.47 65.72 72.31 76.01 78.73 69.84 71.72 76.21 91.35 93.04 93.61
MECT (Wu et al., 2021) 58.55 60.77 66.13 72.82 75.85 79.16 70.54 73.87 76.48 91.52 93.63 93.90
LEBERT (Liu et al., 2021) 61.23 64.03 67.63 72.39 75.00 77.88 71.12 74.46 77.44 93.08 94.16 95.05
PCBERT (Lai et al., 2022) 70.73 70.78 72.81 77.67 81.96 83.66 73.32 75.41 79.21 93.63 94.31 95.18
BiaffineNER (Yu et al., 2020) 58.74 66.41 69.70 77.28 80.21 81.68 74.62 76.98 79.46 93.30 94.61 95.49
[…]
CNN Nested NER (Yan et al., 2023) 64.81 67.75 69.96 79.28 81.58 83.94 75.39 77.86 80.06 93.10 94.54 95.39
DiFiNet (Cai et al., 2024) 67.35 69.02 72.19 79.81 81.32 83.29 75.40 77.05 79.61 93.81 94.75 95.93
KnowFREE + Fusion (GLM3-6B) 66.76 71.59 72.99 79.30 82.13 84.50 76.31 78.55 80.53 94.03 95.04 96.14
[…]
Dataset / Model CMeEE-v2 PKU MSR UD
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003) - - - 93.49 94.31 95.02 90.60 91.73 92.93 87.11 89.87 91.98
FLAT (Li et al., 2020) - - - 78.28 80.10 80.22 77.72 78.27 78.44 77.76 78.39 80.11
MECT (Wu et al., 2021) - - - 87.28 87.54 87.58 87.48 87.53 87.71 87.12 87.30 87.61
LEBERT (Liu et al., 2021) - - - 93.42 94.09 94.97 90.64 91.95 93.27 88.93 91.85 93.43
PCBERT (Lai et al., 2022) - - - 93.47 94.07 94.54 90.01 92.18 93.10 89.81 91.70 93.67
BiaffineNER (Yu et al., 2020) 61.26 65.56 68.40 93.20 94.39 94.94 91.60 92.63 93.64 88.84 90.74 92.68
W2NER (Li et al., 2022) 61.10 65.67 68.72 93.90 94.64 95.41 91.61 92.76 93.82 90.12 92.91 94.86
CNN Nested NER (Yan et al., 2023) 62.32 66.43 69.06 94.00 94.59 95.47 91.72 92.86 93.67 90.21 92.48 94.70
DiFiNet (Cai et al., 2024) 63.55 66.26 67.28 93.81 94.76 95.26 91.46 92.45 93.24 90.94 92.82 94.62
KnowFREE + Fusion (GLM3-6B) 63.64 67.47 69.52 94.07 94.94 95.51 91.73 92.91 93.92 90.99 92.71 95.00
[…]
KnowFREE + Fusion (Qwen-14B) 63.23 67.25 69.19 94.36 94.94 95.51 91.76 92.93 93.89 90.53 93.03 94.97
[…]
Table 1: The overall results in few-shot sequence labeling tasks, where the value highlighted in bold is the best value, and the value underline is the second-best value.
4.1 Experiment Setup
Training: To comprehensively evaluate the effectiveness of our data augmentation strategy on different LLMs, and to fairly compare previous related methods. We use GLM3-6B (GLM et al., 2024), GLM4-9B (GLM et al., 2024) and Qwen2.5-14B (Yang et al., 2024b, a) as LLMs for label extension annotation and enriched explanation synthesis. And we choose BERT (Devlin et al., 2019) as the backbone model of KnowFREE encoder. In the pipeline of label extension annotation, the pre-trained embedding model of $\mathcal{M}$ is set as “text2vec” (Ming, 2022), the Top-$p$ value in Equation 7 is set as five, and the threshold $\epsilon$ is set as 1.5. In the training stage, the hidden size $D$, $D^{\prime}$, and $\tilde{D}$ are set as 768, 200, and 200, respectively. The activation function of $\sigma$ and $\sigma^{\ast}$ are defined as Leaky ReLU and GeLU, respectively. In the local multi-head attention module, the window size $\omega$ is set as three and the number of attention heads $K$ is set as ten. The detail settings of $\alpha$, $\beta$ and other parameter are presented in Appendix G.
Evaluation: To evaluate the performance of our method in low-resource sequence labeling tasks, we conducted experiments on multiple datasets. These include four flat NER datasets: Weibo (Peng and Dredze, 2015), Youku (Jie et al., 2019), Taobao (Jie et al., 2019), and Resume (Zhang and Yang, 2018b); a nested NER dataset, CMeEE-v2 (Zhang et al., 2022); two word segmentation datasets, PKU (Emerson, 2005) and MSR (Emerson, 2005); and a POS tagging dataset, UD (Nivre et al., 2016). Real-world datasets often exhibit imbalanced category distributions, making them ideal for evaluating a model’s generalization ability in authentic low-resource scenarios. In our main experiments, we randomly sampled subsets of varying sizes from the datasets to simulate sequence labeling under low-resource conditions. The training set sample sizes $\mathcal{K}$ were set to 250, 500, and 1000, respectively. The detail statistics of all datasets are presented in Appendix F. We control randomness by ensuring that subsequent samples include previously sampled samples.
Baselines: To provide a fair comparison of the effectiveness of our method, we selected several mainstream approaches for Chinese sequence labeling tasks, including BERT-CRF (Devlin et al., 2019; McCallum and Li, 2003), FLAT (Li et al., 2020), MECT (Wu et al., 2021), and LEBERT (Liu et al., 2021). Additionally, we included PCBERT (Lai et al., 2022) that also leverage extension knowledge enhancement, as well as methods designed for nested entity recognition, including BiaffineNER (Yu et al., 2020), W2NER (Li et al., 2022), CNN Nested NER (Yan et al., 2023), and DiFiNet (Cai et al., 2024). In our analysis, we further compared the performance of vanilla LLMs and LoRA fine-tuned models on sequence labeling tasks. Additionally, we evaluated the performance of various LLM-based data augmentation strategies.
4.2 Main Results
As shown in Table 1, our method consistently outperforms others on average. On the Weibo dataset, our best result surpasses PCBERT by 6.4%. Under the $\mathcal{K}=250$ setting, our approach outperforms the strongest baseline by an average margin of 1.95% across all datasets. Furthermore, incorporating the enriched explanation synthesis pipeline (+ Fusion + Synthetic) improves performance over using only label extension annotation (+ Fusion) when $\mathcal{K}$=250 across all datasets. This effect is particularly evident on the Weibo dataset, where our best GLM3-6B-based “+ Fusion + Synthetic” ($\mathcal{K}$=1000) surpasses the performance of “+ Fusion” ($\mathcal{K}$=1000) by 1.79%. However, as $\mathcal{K}$ increases, the performance gains from “+ Fusion + Synthetic” in most datasets gradually decline, and fall slightly below “+ Fusion”. This suggests that enriched data synthesis is more effective in resource-scarce scenarios. Meanwhile, under the “+ Fusion ($\mathcal{K}$=1000)” setting, our best results achieves an average improvement of 0.52% over the strongest baseline across all NER datasets, further validating the effectiveness of the label extension annotation strategy. Finally, we observe that different LLM backbones yield varying performance across datasets. While larger LLMs contribute more significantly to the improvement of “+ Fusion” in low-resource scenarios, they do not exhibit a clear advantage in data synthesis. This may be due to variations in how different LLMs align with prompt instructions. Since our prompts were designed and validated on GLM3-6B, their effectiveness may differ when applied to other LLMs.
4.3 Analysis
Ablation Studies: To assess the impact of each module and strategy in our method, we conducted ablation experiments by removing individual components and evaluating their effects on model performance. The results are summarized in Table 2. “w/o L.E.A.” refers to the model’s performance without the label extension annotation, relying solely on the vanilla KnowFREE model. This configuration results in a significant performance drop compared to the default method. Nevertheless, it still outperforms W2NER and CNN Nested NER on datasets such as Weibo, Youku, CMeEE-v2, and UD, demonstrating the effectiveness of the local multi-head attention module in capturing and enhancing neighborhood feature interactions. “w/o S.L.” evaluates the impact of removing the synonymous label merging strategy, which results in severe performance degradation. This highlights that even with the inclusion of extension label features, the absence of synonymous label merging leads to confusion when equivalent labels are treated as distinct categories, impairing the model’s judgment. “w/o entity” and “w/o pos,” evaluate the removal of extension entities and POS tagging features, respectively. On NER datasets, the absence of extension entity features causes a more significant performance drop compared to POS tagging, highlighting the critical role of entity-related features in improving the model’s ability to detect entities. The impact of “w/o pos” varies across datasets. On MSR and UD, removing extension POS tagging features leads to improved performance. This may be due to noise introduced by the extension features. Since these features are highly correlated with the target labels, any noise in the extension features could disproportionately affect the model’s performance on the target task.
Method Weibo Youku Taobao Resume CMeEE-v2 PKU MSR UD
Default 72.99 84.50 80.53 96.14 69.52 95.51 93.92 95.00
     w/o L.E.A. 72.32 84.26 79.93 96.02 69.49 93.75 93.67 94.88
     w/o S.L. 72.26 84.19 79.80 95.74 69.18 94.86 93.87 94.85
     w/o entity 72.25 84.37 80.14 95.86 69.08 95.51 93.91 94.97
     w/o pos 72.39 84.45 80.48 96.14 69.26 95.51 93.93 95.01
Table 2: Results of F1 scores in ablation studies, all results are trained on datasets with $\mathcal{K}$=1000. The backbone LLM is GLM3-6B.
Figure 4: Performance comparison with and without enriched explanation synthesis under k-shot sampling.
Impact of Enriched Explanation Synthesis in K-Shot Sampling: To evaluate the impact of enriched explanation synthesis on model performance in low-resource scenarios, we conducted experiments following the n-way k-shot paradigm. Using a greedy sampling strategy, we repeatedly sampled data from flat NER datasets, ensuring that each target label in the k-shot setting was represented by at least k instances. The sampled data was then augmented with GLM3-6B. We compared the model’s performance with and without enriched explanation synthesis, as shown in Figure 4. Here, “w Synthetic” indicates the performance with enriched explanation synthesis, while “w/o Synthetic” reflects the performance without it. The results demonstrate that “w Synthetic” achieves a substantial performance boost over “w/o Synthetic” from the outset. Notably, when k is less than 15, enriched explanation synthesis consistently delivers rapid performance improvements across all datasets. As k increases, the performance gap narrows, but “w Synthetic” continues to outperform “w/o Synthetic” across all settings. These findings highlight that in resource-scarce scenarios, synthesizing enriched data is more effective than directly injecting features into raw samples. These findings highlight the critical role of enriched explanation synthesis in enhancing model performance, particularly when labeled data is limited.
Besides, we conduct extra ablation studies to analyze the effect of the number of heads in local attention in Appendix C. Additionally, we evaluate sequence labeling performance with LLMs in Appendix A and compare various data synthesis strategies in Appendix D. Finally, we visualize the logit scores of extension labels in Appendix E.
5 Conclusion
In this paper, we propose a novel framework that integrates an LLM-based knowledge enhancement workflow with a span-based sequence labeling model. Our approach improves model performance by generating contextual interpretations of target entities and annotating extension labels. Additionally, our KnowFREE model effectively incorporates extension label features to enhance extraction capabilities. Extensive experiments on Chinese few-shot sequence labeling datasets demonstrate that our method achieves state-of-the-art performance, showcasing its effectiveness and efficiency.
Limitations
While enriched explanation synthesis significantly improves model performance in low-resource scenarios (e.g., with fewer than 500 original samples), its effectiveness diminishes as the size of the original dataset increases. Specifically, when the number of original samples exceeds this threshold, distributional discrepancies between synthetic samples and target domain semantics can lead to the synthetic data having a negative impact that outweighs its benefits. In future work, we plan to explore adaptive alignment mechanisms to better align synthetic and original data across different data scales.
Ethics Statement
Our data augmentation method utilizes LLMs to generate data independently of the existing training set. However, the generated data may reflect social biases inherent in the pre-training corpus. To mitigate the risk of propagating biased information into sequence labeling models, we recommend conducting manual reviews before integrating the synthesized data into practical applications.
Appendix A Sequence Labeling with LLMs
Dataset / Model Weibo Youku Taobao Resume CMeEE UD PKU MSR
GLM3-6B 3.28 9.84 9.21 13.71 2.41 1.15 3.76 6.69
GLM4-9B 12.43 31.69 14.90 47.12 16.28 26.38 26.18 31.12
Qwen2.5-14B 31.06 51.00 5.89 51.67 22.38 10.06 2.33 3.02
Table 3: Results on sequence labeling tasks with vanilla LLMs.
Dataset / Model Weibo Youku Taobao Resume
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
GLM3-6B (LoRA) 12.29 20.44 28.22 34.52 38.62 39.29 29.91 30.02 30.03 39.92 40.76 39.98
GLM4-9B (LoRA) 49.20 54.54 60.30 73.79 74.56 74.84 58.84 63.63 69.20 83.31 86.24 88.27
Qwen2.5-14B (LoRA) 50.93 56.87 56.30 72.89 75.10 78.83 61.58 59.43 66.98 84.45 88.12 88.65
Dataset / Model CMeEE-v2 PKU MSR UD
$\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000 $\mathcal{K}$=250 $\mathcal{K}$=500 $\mathcal{K}$=1000
GLM3-6B (LoRA) 12.75 13.30 13.55 20.62 26.27 29.21 23.53 26.04 29.38 32.20 37.57 37.81
GLM4-9B (LoRA) 50.67 54.50 56.22 71.41 73.68 75.02 75.70 78.64 80.94 72.32 77.50 79.69
Qwen2.5-14B (LoRA) 53.59 53.80 59.31 71.94 72.71 75.91 77.83 79.89 80.07 74.40 77.04 80.70
Table 4: Performance of LLM fine-tuning with LoRA on sequence labeling tasks.
To evaluate the performance of directly using LLMs for sequence labeling tasks in low-resource scenarios, we present the results of vanilla LLMs in Table 3 and the performance of LoRA-finetuned models trained on sampled datasets in Table 6. During zero-shot inference, LLMs extract entities based on the input text and target labels. As shown in Table 3, vanilla LLMs exhibit significantly poor performance on all datasets, likely due to their limited understanding of target label definitions. Therefore, for domain-specific sequence labeling tasks, incorporating few-shot examples into the prompt or fine-tuning the model with LoRA could be a more effective and practical approach.
In the experiments with LoRA fine-tuning, all LLMs show performance improvements compared to zero-shot inference. Compared to other LLMs, GLM3-6B still underperforms in LoRA fine-tuning, likely due to its weaker ability to align with the target domain of sequence labeling. As a result, the synthesized data produced by GLM3-6B contains a considerable amount of irrelevant information. With $\mathcal{K}$=250, Qwen2.5-14B demonstrates outperforms other LLMs across all datasets, suggesting that models with larger parameters tend to show enhanced initial performance in low-resource contexts. Nonetheless, with the sample size increases, the performance improvements of Qwen2.5-14B on datasets such as Weibo, Taobao, and MSR fell short compared to GLM4-9B. This could be due to variations in model performance when applied to different domain-specific data distributions. It is important to highlight that the performance of LLMs on most datasets was still below that of conventional sequence labeling baselines, including the the relatively simple BERT-CRF. The findings from the main results indicate that the knowledge contained in LLMs can significantly improve sequence labeling performance. However, they lack the necessary expertise and alignment capabilities for handling domain-specific datasets. Due to biases in domain data distribution, LLMs struggle to identify target entities in particular fields as efficiently as conventional sequence labeling techniques.
Considering the trade-offs between cost and performance, the data augmentation approach leveraging LLMs proposed in this study offers a more practical and efficient solution. This method bridges the gap between LLMs’ general knowledge and the specialized requirements of domain-specific sequence labeling tasks.
Appendix B Visual Analysis of Enriched Explanation Synthesis
Figure 5: t-SNE visualization of the training, test and enriched explanation samples under different sampling sizes. The synthetic enriched explanation samples are generated by GLM3-6B, and they are represented by the “Synthetic” in the legend.
In this section, we use the BERT-based embedding model, text2vec (Ming, 2022), to generate sentence embeddings for the original training and test samples, as well as the enriched explanation samples from the Weibo, Youku, Taobao, Resume, and CMeEE-v2 datasets. These embeddings are projected into a two-dimensional space using the t-SNE algorithm, with the results shown in Figure 5. At $\mathcal{K}$=250, the sparse distribution of training samples in datasets such as Weibo, Youku, Taobao, and Resume fails to fully cover the semantic space of the test sets, the limitation is observed across all datasets. However, the synthesized samples effectively bridge these gaps in the semantic space, leading to significant performance improvements in low-resource scenarios. As $\mathcal{K}$ increases, the training samples begin to provide more comprehensive coverage of the semantic space for most datasets. At $\mathcal{K}$=1000, however, semantic distribution discrepancies are noticeable between some training and synthesized samples in datasets like Youku, Taobao, and CMeEE-v2, potentially introducing noise that may adversely affect model performance. On the Weibo dataset, certain regions of the test set’s semantic space remain underrepresented by the original training samples, even at $\mathcal{K}$=1000. This explains why models trained with synthesized samples continue to exhibit notable performance advantages over those trained without synthesized samples at this sample size.
These observations further underscore the effectiveness of the enriched explanation synthesis method in improving model performance under low-resource conditions. However, they also highlight that as the sample size increases, the potential adverse effects of synthesized samples, such as semantic noise, become more pronounced.
Appendix C Impact of the Number of Heads in Local Attention
Figure 6: Performance comparison between different data synthesis strategies under k-shot sampling.
Figure 7: Performance variation with different numbers of heads in local attention module.
To analyze the impact of different numbers of attention heads on model performance, we conducted experiments on four flat NER datasets: Weibo, Youku, Taobao, and Resume, using a sampling size of 1000. The model was trained on the sampled data with extension labels extracted by GLM3-6B, and the results are presented in Figure 7.
The results show that the number of attention heads has a moderate effect on performance. Overall, as the number of heads increases, the model’s performance improves slightly, though the gains are not particularly significant. Additionally, it is important to consider how feature vectors are distributed across the heads. Based on these findings, we chose to use ten attention heads in this study.
Appendix D Comparison of different data synthesis strategies
To evaluate the effectiveness of the enriched explanation synthesis strategy, we reproduced and compared two LLM-based data synthesis methods designed for sequence labeling tasks: LLM-DA (Ye et al., 2024) and ProgGen (Heng et al., 2024). These methods were used to synthesize data from k-shot samples of the original datasets. For consistency, all synthesized samples were annotated using the KnowFREE model trained on the original data without synthesized samples. We conducted experiments on the Weibo, Youku, Taobao, and Resume datasets, all results are presented in Figure 6. The results show that the performance gains from all data synthesis strategies decrease as the number of samples increases. Notably, in scenarios with $\text{k}\leq 30$ on the Youku and Taobao datasets, both LLM-DA and ProgGen lead to performance degradation compared to models trained without synthesized data. This suggests that the synthesized samples generated by these methods may contain inherent semantic distribution biases, which diminish their effectiveness in enhancing performance in certain low-resource domains.
In contrast, our method consistently delivers significantly better performance improvements for $\text{k}\leq 50$, with particularly notable gains on the Weibo dataset. These results demonstrate that the samples synthesized by our approach are more closely aligned with the target domain’s distribution and exhibit superior robustness.
Appendix E Visualization of the Logits with Extension Labels
Figure 8: Heatmap visualization of logits scores between target labels and extension labels on the test sets.
To further investigate the interaction between the introduced extension labels and target labels in the model, we visualized the logit scores of extension labels corresponding to each predicted target label position in the test set. These scores were aggregated by summing and averaging across label categories, and the results are displayed as a heatmap in Figure 8. In each heatmap subplot, the horizontal axis represents the target labels, while the vertical axis corresponds to the extension labels.
The results reveal that certain extension labels exhibit strong correlations with specific target labels. For instance, in the Weibo dataset, both “PER.NAM” and “PER.NOM” show a notable association with the extension label “PERSON”. Similarly, in the Resume dataset, “COUNT” demonstrates strong correlations with the extension labels “Country,” “Location,” and “description”. Incorporating these significant relationships during training allows the model to leverage co-occurrence patterns, enhancing its ability to perform fine-grained semantic understanding and improving target entity prediction.
However, some extension labels were observed to have strong correlations with all target labels. Since the KnowFREE model produces independent probability scores for each target label, the influence of such extension labels on target entity predictions is generally limited when their number is small, as their training weights are relatively low. Conversely, when these extension labels become overly numerous, they can negatively impact model training. In such cases, reducing their weights further can help mitigate these effects and enhance overall model performance.
Appendix F Statistics of Datasets
Dataset Dev Test Label Types Domain
Weibo 0.27k 0.27k 8 Social Media
Youku 1.00k 1.00k 3 Video Content
Taobao 1.00k 1.00k 4 E-commerce
Resume 0.46k 0.48k 8 Human Resources
CMeEE-v2 4.98k 4.98k 9 Medical
PKU 1.00k 2.04k 1 News
MSR 1.00k 3.99k 1 News
UD 0.50k 0.50k 16 News, Literature
Table 5: Statistics of development sets, test sets, label types and domains of all datasets.
The detailed statistics of the datasets are shown in Table 5. These datasets span various domains, including Social Media, E-commerce, and Medical, enabling a comprehensive evaluation of the model’s performance across different fields.
Appendix G More Experiment Settings
In this section, we describe the additional experimental parameter settings for our method. In the sequence labeling model, we distinguish between the learning rate for the PLM and other modules, setting them to 2e-5 and 1e-3, respectively. For the weight $\alpha$ of extension labels, we introduce a dynamic weight calculation mechanism to handle the influence of frequently occurring extension labels (e.g., POS tags). These frequent labels can affect the gradient calculation, leading to reduced attention to target labels. To address this, we calculate the count $C_{i}$ of each extension label and the average count $\hat{C}$ of target labels, and then compute the weight coefficient $\alpha_{i}$ as follows:
$\alpha_{i}=0.5\times(\hat{C}/C_{i}).$ (20)
Dataset Weight Decay $\beta$ Type
Weibo 1e-3 1.0 Flat NER
Youku 1e-2 0.4 Flat NER
Taobao 1e-3 0.4 Flat NER
Resume 1e-3 0.4 Flat NER
CMeEEv2 1e-3 0.4 Nested NER
PKU 1e-2 1.0 Word Segment
MSR 1e-2 1.0 Word Segment
UD 1e-2 1.0 POS Tagging
Table 6: Settings of $\beta$ and weight decay across different datasets.
The training weight for synthesized samples ($\beta$) and the weight decay parameter are provided in Table 6. As shown, for most NER datasets with more complex entity semantics, we use a smaller weight decay parameter to improve model fitting during training. In contrast, for POS tagging and tokenization datasets, we apply a larger weight decay to prevent overfitting. Additionally, for NER datasets, where entity labels are more prone to noise from synthesized samples, we set $\beta=0.4$. On the other hand, for datasets with strong baseline performance, increasing $\beta$ to 1.0 helps the model better utilize synthesized samples during training. Our implementation is built on the Huggingface Transformers (Wolf et al., 2020), and all experiments are conducted using two NVIDIA A6000 GPUs for both training and inference.
Appendix H Prompts
In this section, we present detailed examples of our workflow prompts for label extension annotation in Figure 10, 10 and enriched explanation synthesis in Figure 12, 12. Since the target dataset is entirely in Chinese, all original prompts are written in Chinese. The English portions in the prompt examples are translations of the original prompts.
Figure 9: The entity extraction prompt in label extension annotation.
Figure 10: The POS tag extraction prompt in label extension annotation.
Figure 11: The entity explanation prompt in enriched explanation synthesis.
Figure 12: The extension description prompt in enriched explanation synthesis.